A triangle sail has an area of 76 square feet. Find the base and the height if the height measure is three more than twice the base. (use formula A= 1/2 bh)
Answer by mananth(16949)

base == b
h= 3+2b
76 = 1/2 *b*(3+2b)
152= 3b+2b^2
2b^2+3b-152=0
2b^2+19b-16b-152=0
2b(b+9.5)-16(b+9.5)=0
(b+9.5)(2b-16)=0
2(b+9.5)(b-8)=0
so b= 8 which is positive.
base = 8 ft
and height = 3+2b= 19 ft